# Pychop: Emulating Low-Precision Arithmetic in Numerical Methods and Neural Networks

### v2

@@ 1. Introduction
The advent of hardware architectures that natively support low-precision arithmetic has catalyzed a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical linear algebra and deep neural network (DNN) training. These algorithms, which strategically integrate low-precision and high-precision computations, have emerged as a focus of research due to their capacity to optimize algorithmic performance across various areas—most notably energy efficiency, computational speedup, and resource utilization—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022). This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling framework for addressing the escalating demands of large-scale scientific simulations and machine learning applications. The ability to tailor precision to specific computational tasks not only reduces power consumption but also reduces processing times, making mixed-precision techniques indispensable in the era of exascale computing and resource-constrained environments.
@@
The advance of mixed-precision training has been a milestone for deep learning efficiency. Micikevicius et al. (2018) pioneered its practical application, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs. This approach has been standardized in PyTorch (Paszke et al., 2019) via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates precision switching and gradient scaling, reducing memory usage by approximately 50% for models like ResNet-50 (He et al., 2016). However, AMP’s dependence on hardware-supported fp16 limits its adaptability to other formats, a constraint our emulator eliminates by supporting arbitrary precision configurations.
Central to this research domain is the advancement of quantization methodologies, encompassing binary, ternary, and 4- to 8-bit schemes, which facilitate substantial compression of model parameters and intermediate representations with minimal degradation in performance (Jacob et al., 2018). Such techniques are pivotal for reducing memory footprints and computational overhead, rendering them especially valuable for deploying DNNs on edge devices and other hardware with limited resources. To preserve the integrity of the compressed information, it is essential to devise quantization strategies that explicitly maintain similarity between the original and quantized representations, a challenge that necessitates sophisticated similarity-preserving algorithms (Zhu et al., 2021). Nevertheless, the deployment of large language models (LLMs) at scale introduces additional complexities, as achieving high availability and scalability remains a significant hurdle. Overcoming these obstacles requires innovative solutions in lexical search, data re-ranking, and retrieval-augmented generation, all optimized to function efficiently at reduced computational costs (Lewis et al., 2020). The successful integration of these techniques not only enhances the efficacy of LLMs but also broadens their applicability, enabling real-time inference and supporting data-intensive research endeavors across diverse domains. As the field progresses, the synergy between mixed-precision computing and quantization promises to redefine the boundaries of computational efficiency and model performance, paving the way for more sustainable and accessible artificial intelligence systems.
@@
Table 1. Key parameters of six floating point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
@@
Our work addresses this limitation with a mixed-precision emulation software module for Python, offering a highly configurable platform to simulate arbitrary low precision floating-point formats and a diverse array of rounding modes. This emulator transcends the constraints of fixed hardware by allowing users to define custom precision configurations—specifying exponent and significand bits—and to select from rounding modes: round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability). Its advantages are substantial and multifaceted:
  • Unmatched Flexibility: Emulates standard formats (see Table 1) and custom designs, enabling researchers to prototype hypothetical hardware or explore theoretical precision limits without physical constraints.
  • Precision Granularity: Provides precise control over numerical representation, critical for dissecting the impact of quantization on gradient updates, weight distributions, and activation ranges in neural networks.
  • Seamless Integration: Embeds directly into PyTorch layers, allowing practitioners to experiment with mixed precision in production-grade training pipelines with minimal overhead.
  • Rounding Mode Exploration: Supports a comprehensive set of rounding strategies, facilitating detailed studies of their effects on numerical stability, convergence rates, and final model performance—an underrepresented area in hardware-based implementations.
This paper is organized as follows: Section 2 reviews prior work on precision emulation software, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance and its value in advancing mixed-precision emulation for deep learning. Section 5 concludes the paper and outlines future work.
@@ 2. Related Work
GNU MPFR (Fousse et al., 2007) C custom A A O $\checkmark$ ! $\checkmark$ $\checkmark$
@@
The columns are categorized as follows: (i) Package name, Primary language, Storage format: The name of the package, its primary programming language, and the supported storage formats. (ii) Target format parameters: $p$: Number of bits of precision in the significand—arbitrary (A) or restricted to the storage format’s significand (R); $e$: Exponent range—arbitrary (A), a sub-range of the storage format (S), or a sub-range only for built-in types (B); $s$: Support for subnormal numbers—supported (Y), not supported (N), supported only for built-in types (B), supported by default but can be disabled (F), or not supported by default but can be enabled (O); built-in: Floating-point formats natively built into the system (e.g., fp16, bf16, tf32). (iii) Rounding modes: Supported modes include round-to-nearest with ties-to-even (RNE), ties-to-zero (RNZ), ties-to-away (RNA), round-toward-zero (RZ), round up/down (RUD), round-to-odd (RO), and stochastic rounding variants (SR). A $\checkmark$ indicates full support, while ! denotes modes that can be simulated at a higher computational cost. (iv) FPQ: support fixed-point quantization. (v) IQ: support integer quantization. (vi) NN: support neural network quantization. (vii) STE: support Straight-Through Estimator, which permits gradients to propagate through during the backward pass that enables the continuation of the backpropagation algorithm.
@@
Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019). Several software libraries have been developed to simulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases. In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
@@
GNU MPFR (Fousse et al., 2007) is a C library designed for multiple-precision floating-point computations with guaranteed correct rounding. It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$), making it a preferred choice for applications demanding high numerical accuracy, such as symbolic computation and numerical analysis. It supports rounding modes including round-to-nearest with ties-to-even (RNE), round-toward-zero (RZ), and round up/down (RUD). However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows. Additionally, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its applicability in machine learning contexts.
@@
SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding. It supports multiple storage formats, restricted significand precision ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$). SIPE implements RNE and RZ rounding modes, ensuring numerical stability for low-precision simulations, such as those in embedded systems. Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility. Moreover, SIPE is not designed for neural network training, making it more suitable for general numerical simulations rather than machine learning applications.
@@
FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020), both C++ libraries, provide frameworks for customized floating-point arithmetic in low-precision simulations. They support restricted significand ($p$) and exponent range ($e$), fully support subnormal numbers ($s$), and use fp32/fp64 storage formats for compatibility with standard representations. Their simplicity, with support limited to RNE rounding, makes them accessible for educational purposes and prototyping. However, this restricted rounding support hampers their adaptability, and neither library includes built-in formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation. FloatX and FlexFloat is constrained by adhering strictly to the ”natural” C++ rules, such as the ”round-to-nearest, ties-to-even” rounding mode and standard datatype casting conventions. While these rules ensure consistency with native floating-point behavior, they may limit users who wish to explore a broader range of rounding strategies across various numerical simulation domains. This restriction can be a drawback for researchers and practitioners seeking more flexibility in modeling custom rounding behaviors tailored to specific applications.
@@
CPFloat (Fasi and Mikaitis, 2023), a C library, is optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), with flexible subnormal number support (F). It includes built-in fp16, bf16, and tf32 formats, enhancing compatibility with hardware-accelerated systems such as GPUs. CPFloat’s comprehensive support for rounding modes—RNE, RNZ, RNA, RZ, RUD, RO, and SR—makes it highly versatile for numerical simulations requiring diverse rounding strategies. However, CPFloat is not designed for neural network training, focusing instead on general-purpose low-precision arithmetic, such as in algorithm development and hardware emulation.
QPyTorch (Zhang et al., 2019) is specifically designed for low-precision arithmetic for neural network training. QPyTorch, a PyTorch-integrated Python tool, uses fp32 storage with restricted $p$ and $e$, lacks $s$ support, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused kernels. However, its limited rounding support (lacking RZ, RUD, RO) restricts advanced quantization studies and numerical simulations.
All aforementioned libraries, except QPyTorch, lack Straight-Through Estimator (STE) support and are thus restricted to post-training quantization (PTQ), rendering them ineffective for quantization-aware training (QAT). Our Pychop library supports fp32/FP64 storage and built-in fp16 and bf16 formats, constrained precision ($p$) and exponent ($e$), and flexible significand ($s$), with rounding modes (RNE, RZ, RUD, RO, SR). It automatically detects tensor gradient information to enable STE, enhancing quantization research flexibility. Pychop integrates seamlessly with NumPy, JAX, and PyTorch—outperforming QPyTorch’s PyTorch-only scope—and offers efficient rounding. Its versatility and multi-framework compatibility make Pychop a superior tool for quantized neural network training across diverse deep learning workflows and scientific computations.
@@ 3. The Pychop library
Real numbers in digital systems must be approximated using discrete representations due to finite storage constraints. Two fundamental approaches dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments. This section provides rigorous definitions for both methods, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks. Our mixed-precision emulation software is implemented as a modular Python code within NumPy, PyTorch, and JAX, comprising three primary components: the Chop/Chopf/Chopi class for core quantization logic and various components in the layers class for neural network integration, supporting various roundings (see Table 3 for details) and subnormal numbers. Here, we detail its design and implementation, emphasizing its main principles and user-friendly features.
@@ 3.2. Implementation details
To round to low precision binary floating point numbers, Pychop provides two modules, namely Chop and LightChop. Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization. Different backends of LightChop offer different features. The Torch and JAX backend offers GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
@@
Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating point emulation features. We demonstrate their basic usage below:

### v4

@@ 1. Introduction
The increasing support of low-precision arithmetic in hardware architectures has triggered a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical analysis and deep learning. There are numerous benefits to using numerical formats with lower precision than single precision primarily because they require less memory bandwidth. The mixed-precision algorithms, which strategically combine low-precision and high-precision computations, have emerged as a heating topic of research due to their capability to enhance algorithmic performance across various domains—most notably energy efficiency, data transfer, and arithmetic speedup—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022). This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling technique for addressing the escalating demands of large-scale numerical simulations and machine learning applications. The ability to tailor precision to specific computational tasks not only reduces energy consumption but also reduces processing times, making mixed-precision computations indispensable in the era of exascale computing and resource-constrained environments.
@@
It is known that the increasing size of neural networks typically improves model generalization ability and accuracy at the cost of memory and compute resources for model deployment. The advance of mixed-precision training has been a milestone for deep learning efficiency. Micikevicius et al. (2018) established its use in practice, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs. This approach has been integrated into PyTorch (Paszke et al., 2019) and implemented via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates training in low precision with gradient scaling, lowering memory consumption by approximately 50% for models like ResNet-50 (He et al., 2016). However, AMP’s dependence on hardware-supported fp16 hinders its flexibility to other floating-point formats, a constraint our emulator eliminates by supporting arbitrary precision configurations. Further, (Perez et al., 2023) propose the training method via the scalings for fp8 (8-bit floating-point with two formats (Noune et al., 2022; Micikevicius et al., 2022)—exponent bits and 2 significand bits for E5 and 4 exponent bits and 3 significand bits for E4) linear layers by dynamically updating per-tensor scales for the weights, gradients and activations, validating an acceptable performance of GPT and Llama 2 deployed with fp8 precision across model sizes ranging from  [^1] M to 70B.
Central to this research domain is the advance of integer quantization methodologies, including binary, ternary, and 4- to 8-bit schemes (Xi et al., 2023, 2024), which significantly compress model parameters and intermediate representations while maintaining performance (Jacob et al., 2018). (Xi et al., 2023) introduce a training method for transformers with all matrix multiplications implemented with the 4-bit integer arithmetic. These methods are crucial for minimizing memory usage and computational demands, rendering them especially valuable for deploying DNNs on edge devices and other hardware with limited resources. To make the compressed information useful, it is critical to devise quantization strategies that explicitly maintain similarity between the original and quantized representations, a challenge that necessitates sophisticated similarity-preserving algorithms (Zhu et al., 2021) that enables more efficient training and deployment of larger neural networks.
@@
Table 1. Key parameters of six floating-point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
@@
We design a low-precision emulation software module called pychop for Python, offering a highly customizable setting for users to simulate arbitrary low-precision arithmetic—floating-point, fixed-point, and integer—with a diverse of rounding modes. This emulator transcends the constraints of fixed hardware by allowing users to define custom precision formats—specifying exponent and significand bits—and to select from rounding modes, e.g., round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability). The software promotes comprehensive precision format support, flexibility, and a beginner-friendly interface:
  • Unmatched Flexibility: Emulating both standard (see Table 1) and customized numerical formats for low-precision emulation, enabling researchers to prototype hypothetical hardware or explore theoretical precision boundaries without physical constraints.
  • Precision Granularity: Providing precise control over numerical representation, which is critical for dissecting the impact of quantization on both numerical algorithms and deep learning methods, e.g., gradient updates, weight distributions, and activation ranges in neural networks.
  • Seamless Integration: Offering direct emulators into PyTorch layers, enabling practitioners to experiment with mixed precision in training and inference pipelines with minimal overhead. The user-friendly API enables users to deploy quantization-aware and post-quantization strategies easily.
  • Rounding Mode Exploration: Supporting a diverse collection of rounding modes is available, enabling empirical evaluation of their roles in numerical stability.
This paper is organized as follows: Section 2 reviews prior work on precision emulation software Pychop, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance in MATLAB and Python and its value in advancing mixed-precision emulation for deep learning applications, namely image classification and object detection. Section 5 concludes the paper and outlines future work.
@@ 2. Related Work
GNU MPFR (Fousse et al., 2007) C custom A A O $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
The columns are categorized as follows: (i) Package name, Primary language, Storage format: The name of the package, its primary programming language, and the supported storage formats. (ii) Target format parameters: $p$: Number of bits of precision in the significand—arbitrary (A) or restricted to the storage format’s significand (R); $e$: Exponent range—arbitrary (A), a sub-range of the storage format (S), or a sub-range only for built-in types (B); $s$: Support for subnormal numbers—supported (Y), not supported (N), supported only for built-in types (B), supported by default but can be disabled (F), or not supported by default but can be enabled (O); built-in: Floating-point formats natively built into the system (e.g., fp16, bf16, tf32). (iii) Rounding modes: Supported modes include round-to-nearest with ties-to-even (RNE), ties-to-zero (RNZ), ties-to-away (RNA), round-toward-zero (RZ), round up/down (RUD), round-to-odd (RO), and stochastic rounding variants (SR). A $\checkmark$ indicates full support. (iv) FPQ: support fixed-point quantization. (v) IQ: support integer quantization. (vi) NN: support neural network quantization. (vii) STE: support Straight-Through Estimator, which permits gradients to propagate through during the backward pass that enables the continuation of the backpropagation algorithm.
@@
Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019). Several software libraries have been developed to emulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases. In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
@@
GNU MPFR (Fousse et al., 2007) is a C library for simulating multiple-precision floating-point computations with guaranteed correct rounding. It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$). It is a preferred choice for numerical simulations demanding high accuracy, such as symbolic computation and numerical analysis. However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows. Besides, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its capability in low-precision emulation in deep learning.
@@
SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding. It supports multiple storage formats, restricted significand ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$). SIPE implements RNE and RZ rounding modes for low-precision simulations which is suitable for exploring numerical stability and precision trade-offs in low-bitwidth computations, such as those encountered in embedded systems. Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility.
@@
FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020) are C++ libraries that provide frameworks for customized floating-point arithmetic in low-precision simulations. These libraries support restricted significand ((p)) and exponent range ((e)), fully accommodate subnormal numbers ((s)), and utilize fp32/fp64 storage formats for compatibility with standard representations. Their simplicity, limited to round-to-nearest-even (RNE) rounding, makes them accessible for educational purposes and prototyping. However, this restricted rounding support limits their adaptability, and neither library includes predefined formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation. Both FloatX and FlexFloat strictly follow standard C++ conventions, including round-to-nearest, ties-to-even rounding, and default datatype casting. Although these conventions ensure compatibility with native floating-point operations, they limit users’ ability to explore diverse rounding strategies in numerical simulations. This constraint poses a notable challenge for researchers and practitioners needing flexible, application-specific rounding behaviors.
@@
CPFloat (Fasi and Mikaitis, 2023) is a C library optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage with restricted significand ($p$), and exponent range ($e$). It includes built-in fp16, bf16, and tf32 formats with flexible subnormal number support (F), enhancing compatibility with hardware-accelerated systems such as GPUs. CPFloat supports a comprehensive set of rounding modes that makes it highly versatile for observing the numerical behavior of low-precision arithmetic. However, CPFloat is not designed for neural network quantization but rather for general-purpose low-precision arithmetic in numerical algorithms.
QPyTorch (Zhang et al., 2019) is a PyTorch-based low-precision simulator, specifically designed for low-precision arithmetic for neural network training without relying on low-precision hardware. As mentioned in (Fasi and Mikaitis, 2023), its principles are analogous to those of chop, in that numbers are stored in binary32 before as well as after rounding, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused CUDA kernels. However, infinities, NaNs, and subnormals are excluded for efficiency because neural network training typically doesn’t involve these special values. Besides, its limited rounding support (lacking RZ, RUD, RO) restricts quantization studies and numerical simulations.
All aforementioned libraries, except QPyTorch, lack Straight-Through Estimator (STE) support and are thus restricted to post-training quantization (PTQ), rendering them ineffective for quantization-aware training (QAT). Our Pychop library supports fp32/FP64 storage and built-in fp16 and bf16 formats, constrained precision ($p$) and exponent ($e$), and flexible significand ($s$), with rounding modes (RNE, RZ, RUD, RO, SR). It automatically detects tensor gradient information to enable STE, enhancing quantization research flexibility. Pychop integrates seamlessly with NumPy, JAX, and PyTorch—outperforming QPyTorch’s PyTorch-only scope—and offers efficient rounding. Its versatility and multi-framework compatibility make Pychop a superior tool for quantized neural network training across diverse deep-learning workflows and scientific computations.
@@ 3. The Pychop library
Numbers in machines are often approximated with truncation using discrete representations due to the constraints arisen from finite storage. Two fundamental representations dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments. This section formulates both representations, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks. Our mixed-precision emulation software is implemented as a modular Python code supported by the backends of NumPy, PyTorch, and JAX, and it comprises three primary components: the Chop, Chopf, and Chopi class for float point, fixed point, and integer quantization, respectively, which are also integrated in the layers class with straight-through estimator for neural network deployment, supporting emulation for various representation format (see Table 3 for details) as well as subnormal numbers. In the following, we detail its design and implementation, emphasizing its main principles and user-friendly features.
@@ 3.2. Implementation details
To round to low precision binary floating-point format, Pychop provides two modules, namely Chop and LightChop. Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization. Different backends of LightChop offer different features. The Torch and JAX backend offers GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
@@
Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating-point emulation features. We demonstrate their basic usage below: